# Model agnostic local variable importance for locally dependent relationships

### v1

@@ 4 Real Data Experiments
4.1 Lichen Classification.
We consider the lichen data where each observation is a different location in a survey region of the Pacific Northwest [7]. This dataset consists of many climate, vegetation, and topographic features for assessing a binary response of the presence or absence of a species of lichen (Lobaria oregana).
We build a Random Forest model using all available features, and extract LIME, SHAP, CLIP, and CLIQUE values. Results for permutation-based global variable importances [5] are shown in Figure 8. MinTempAve (minimum temperature) and ACONIF (age of conifers) are two of the most important variables for classifying the presence of Lichen. Descriptive analyses show that MinTempAve is highly collinear with AmbVapPressAve ($\rho=0.997$) and Elevation ($\rho=-0.978$), while being mostly orthogonal to ACONIF ($\rho=-0.116$). We focus on MinTempAve, recognizing that very similar patterns will exist for AmbVapPressAve and Elevation. The ACONIF PDP shows a steady upward trend, whereas MinTempAve exhibits a plateau, followed by a steep increase, and another plateau. We aim to determine whether the importance of ACONIF changes for different minimum temperatures.
@@
To do this, we split our data based on whether MinTempAve was below 65. The resulting boxplots in Figure 9 show that the value of the ACONIF variable has a much greater CLIQUE importance when minimum temperatures are higher. This makes sense, as lichen struggle to survive in areas with low or freezing temperatures, making ACONIF less important in those regions. If it is sufficiently warm for lichen to survive, then ACONIF influences their success greatly. The SHAP value distributions in Figure 9 show apparent differences in spread when MinTempAve is above or below 65, but the effect is more evident in the CLIQUE comparisons. Aggregate absolute SHAP values yield an impressive ratio of 3.52, while CLIQUE values yield a more outstanding aggregate ratio of 9.58 or an aggregate absolute value ratio of 11.25.
@@
We now demonstrate the use of CLIQUE on individual data points, with a focus on the first two observations. From Figure 10, we see that the CLIQUE value of PctVegCov has the strongest implications for observation 1, while MinTempAve and ACONIF do not appear among the top variables. Observation 2 aligns more with the global story, having MinTempAve as the top variable and ACONIF on the list. The SHAP values somewhat agree with these, although they prioritize strong global variables a bit more. LIME heavily emphasizes global variables here. SHAP values appear to blend the CLIQUE values of Figure 10 and the global importances from Figure 8.
4.2 MNIST Digit Classification.
As another illustrative example, we analyze a down-sampled version of the MNIST Digit test dataset [1] available from scikit-learn [21]. This dataset consists of 64 features, each a pixel value, with x and y-coordinates ranging from 1 to 8. These pixel values help determine which hand-drawn digit each observation corresponds to. Since this is a multi-class problem, SHAP and LIME are not natively adapted to provide local importances. An example plot of digits is show in Figure 11. We train a Random Forest model on this data to obtain CLIQUE values. Results for the top permutation-based global importance variables and PDPs are shown in Figure 12.
@@
The boxplots in Figure 13 show that x4y6 has a greater importance when x3y6 is less than $7.5$. x3y6 has a greater importance when x4y6 is less than $4.5$. These results imply an interaction between these pixels, where each pixel value is more important when the other has a lower value. This behavior is similar to results seen in the AND Gate data of Section 3.1.
@@
In addition to comparing pixel importances against pixel values, we also compare them across digit labels. We present this from two different perspectives. The first is found in Figure 14 which shows CLIQUE boxplots grouped by labels. Pixel x4y6 has noticeably higher CLIQUE values for the digits 3 and 9. Although initially surprising, further analysis reveals that for digits 3 and 9, x4y6 equals 0 more than 80% of the time, while for other digits, it equals 0 less than 60% of the time. Therefore, a value of 0 strongly indicates that a digit is likely a 3 or 9, whereas a value above 0 suggests otherwise.
@@ 5 Discussion and Conclusion
In comparing CLIQUE to SHAP and LIME, we briefly note that the computation time for CLIQUE is roughly similar to both, with SHAP being faster and LIME slower. In assessing the local importances, we observe that LIME values consistently tend to capture marginal information exclusively. This is not surprising, since LIME tends to be focused on explaining individual predictions, and is rarely applied globally as we have done here. SHAP offers a combination of marginal and conditional information, mostly showing the marginal trend, but also sometimes splitting out conditional groups. SHAP also preserves directions and magnitudes, providing immediate insights about how a variable value contributes to a prediction. Positive values imply an increase to an average prediction, while negative values imply a decrease. Global assessments for SHAP are done on the spread of the values, not the center, because the centers are zero.
CLIQUE is highly conditional and focuses on how a variable value contributes to prediction error rather than the prediction itself. Near-zero CLIQUE values indicate that a particular variable has no effect on an individual error, while positive values suggest the variable is important for reducing that error. Global assessments for CLIQUE can be observed at the centers and spreads of values. We note that if a variable is more important in one location, the center and the spread of the CLIQUE values will be higher in that location.
@@
A Additional Experiments
Here we include additional results from a regression dataset for predicting Concrete Strength [29]. We also show extended results from the MNIST digit classification dataset.
A.1 Concrete Regression.
@@ A Additional Experiments
A.2 Additional MNIST Results.
@@
CLIQUE is useful for identifying pixels that help distinguish these digit 1 samples from other digits. While samples A and B share a value of 0 for x1y1, it is not important because most digits also have a value of 0. In contrast, x4y3 has a much higher value for digit 1 compared to most other digits. Both samples A and B have high values for x4y3, making it locally important. The digit 1 also shows much higher values for x5y3 compared to other digits. For sample B, this digit is very important due to its high value. However, sample A has a much lower value for x5y3, making x5y3 much less important for assisting in the classification of sample A.

### v2

@@ 4 Real Data Experiments
4.1 Lichen Classification
In the lichen data, each observation is a different location in a survey region of the Pacific Northwest. This dataset consists of many climate, vegetation, and topographic features for assessing a binary response of the presence or absence of a species of lichen (Lobaria oregana). We build a Random Forest model using all available features, and extract LIME, SHAP, CLIP, and CLIQUE values. Results for permutation-based global variable importances [4] are shown in Figure 8. MinTempAve (minimum temperature) and ACONIF (age of conifers) are two of the most important variables for classifying the presence of Lichen. Descriptive analyses show that MinTempAve is highly collinear with AmbVapPressAve ($\rho=0.997$) and Elevation ($\rho=-0.978$), while being mostly orthogonal to ACONIF ($\rho=-0.116$). We focus on MinTempAve, recognizing that very similar patterns will exist for AmbVapPressAve and Elevation. The ACONIF PDP shows a steady upward trend, whereas MinTempAve exhibits a plateau, followed by a steep increase, and another plateau. We aim to determine whether the importance of ACONIF changes for different minimum temperatures.
@@ 4.1 Lichen Classification
To do this, we split our data based on if MinTempAve was below 65. The resulting boxplots in Figure 9 show that the value of the ACONIF variable has a much greater CLIQUE importance when minimum temperatures are higher. This makes sense, as lichen struggle to survive in areas with low or freezing temperatures, making ACONIF less important in those regions. If it is sufficiently warm for lichen to survive, then ACONIF influences their success greatly. The SHAP value distributions in Figure 9 show apparent differences in spread when MinTempAve is above or below 65, but the effect is more evident in the CLIQUE comparisons. Aggregate absolute SHAP values yield an impressive ratio of 3.52, while CLIQUE values yield a more outstanding aggregate absolute value ratio of 11.25.
@@
We now demonstrate the use of CLIQUE on individual data points, with a focus on the first two observations. From Figure 10, we see that the CLIQUE value of PctVegCov is the most important for observation 1, while MinTempAve and ACONIF do not appear among the top variables. Observation 2 aligns more with the global story, having MinTempAve as the top variable and ACONIF on the list. LIME heavily emphasizes global variables here. SHAP values appear to blend the CLIQUE values of Figure 10 and the global importances from Figure 8.
4.2 MNIST Digit Classification
As another illustrative example, we analyze a down-sampled version of the MNIST Digit test dataset [1] available from scikit-learn [21]. This dataset consists of 64 features, each a pixel value, with x and y-coordinates ranging from 1 to 8. These pixel values help determine which hand-drawn digit each observation corresponds to. Since this is a multi-class problem, SHAP and LIME are not natively adapted to provide local importances. An example plot of digits is shown in Figure 11. We train a Random Forest model on this data to obtain CLIQUE values. Results for the top permutation-based global importance variables and PDPs are shown in Figure 12.
@@ 4.2 MNIST Digit Classification
The boxplots in Figure 13 show that x4y6 has a greater importance when x3y6 is less than $7.5$, while x3y6 has a greater importance when x4y6 is less than $4.5$. These results imply an interaction between these pixels, where each pixel value is more important when the other has a lower value. This behavior is similar to results seen in the AND Gate data of Section 3.1.
@@
In addition to comparing pixel importances against pixel values, we also compare them across digit labels. We present this from two different perspectives. The first is found in Figure 14 which shows CLIQUE boxplots grouped by labels, effectively illustrating that CLIQUE satisfies property P2. Pixel x4y6 has noticeably higher CLIQUE values for the digits 3 and 9. Although initially surprising, further analysis reveals that for digits 3 and 9, x4y6 equals 0 more than 80% of the time, while for other digits, it equals 0 less than 60% of the time. Therefore, a value of 0 strongly indicates that a digit is likely a 3 or 9, whereas a value above 0 suggests otherwise.
@@ 5 Discussion and Conclusion
While this paper focuses on comparing the values themselves, we acknowledge that computation speed is often a key consideration. Time-cost comparison plots and analysis can be found in Appendix A.4. From this analysis we found that SHAP and CLIQUE are noticeably faster than LIME. Additionally, CLIQUE scales linearly with more observations, which is better than the quadratic scaling found in SHAP. Conversely, SHAP scales roughly constant with respect to the number of feature, which is superior to the linear scaling of CLIQUE. The current version of CLIQUE is competitive with SHAP and LIME, with room for further improvements.
In assessing the local importances, we observe that LIME values consistently tend to capture marginal information exclusively. This is not surprising, since LIME tends to be focused on explaining individual predictions, and is rarely applied globally as we have done here. SHAP offers a combination of marginal and conditional information, mostly showing the marginal trend, but also sometimes splitting out conditional groups. SHAP also preserves directions and magnitudes, providing immediate insights about how a variable value contributes to a prediction. Positive values imply an increase to an average prediction, while negative values imply a decrease. Global assessments for SHAP are done on the spread of the values, not the center, because the centers are zero.
CLIQUE is highly conditional, measuring how a variable value affects prediction error rather than the prediction itself. Near-zero values indicate no effect on individual error, while positive values suggest the variable is important for reducing it. Global assessments can potentially be made from the centers and spreads of CLIQUE values. We note that if a variable is more important in one location, the center and the spread of the CLIQUE values will be higher in that location.
@@
Appendix A Additional Experiments
Here we include additional results from a regression dataset for predicting Concrete Strength [28]. We also show extended results from the MNIST digit classification dataset, analyze the effect of $M$, and compare the computational cost of CLIQUE with SHAP and LIME.
A.1 Concrete Regression
@@
A.2 Additional MNIST Results
@@
CLIQUE is useful for identifying pixels that help distinguish these digit 1 samples from other digits. While samples A and B share a value of 0 for x1y1, it is not important because most digits also have a value of 0. In contrast, x4y3 has a much higher value for digit 1 compared to most other digits. Both samples A and B have high values for x4y3, making it locally important. The digit 1 also shows much higher values for x5y3 compared to other digits. For sample B, this digit is very important due to its high value. However, sample A has a much lower value for x5y3, making x5y3 much less important for assisting in its classification.
A.3 Analyzing the effect of M
Figure 21: Plot showing panels of data scatterplots and the relationship between importance value distributions and the hyper-parameter $M$. Top-Left: Scatter-plot showing $V1$ and $V2$, colored by the response ($y$). Top-Right: Scatter-plot showing $V1$ and $V2$, colored by regions of interest with respect to $V1$. Bottom-Left: Plot showing the distribution of CLIP values across the hyper-parameter $M$, colored by regions. Bottom-Right: Plot showing the distribution of CLIQUE values across the hyper-parameter $M$, colored by regions. The CLIQUE values have a smaller spread for each value of $M$.
We used the data structure from Section 3.1 to examine the effect of $M$ on CLIQUE values. Recall that $M$ represents the number of altered values in the quantile grid. To demonstrate some of the advantages that CLIQUE has over permutation-based methods to assessing local importance, we also run the same experiments using CLIP, where $M$ represents the number of permuted values used to assess the importance.
The results are shown in Figure 21. Consistent with our previous experiments, these results show that the CLIP values have a much larger spread than the CLIQUE values in each region of interest.
The results in Figure 21 also provide guidance on selecting an appropriate $M$. As noted in Section 3, we chose $M=25$ to stabilize estimates and reduce variance. Figure 21 reveals that CLIQUE value variances remain remarkably consistent across all $M$, while the value centers show some initial volatility before stabilizing around $M=15$. Beyond this point, higher $M$ values provide diminishing returns at an increased computational cost.
Considering these findings alongside the stability observed in Section 3, we suggest that $M$ values between 20 and 50 are generally conservative enough to yield stable estimates. However, larger $M$ values may be warranted for datasets with highly complex interactions or decision boundaries.
A.4 Time Cost Comparison
Here we performed a computational time comparison between CLIQUE, SHAP, and LIME. We used the MNIST dataset with images aggregated to a 6x6 pixel resolution [23] to test the effects of increasing both the number of observations and the number of features on computational time. To ensure compatibility with both SHAP and LIME (which are adapted natively only to binary classification), we limited the dataset to just two digit classes: 3 and 8. These comparisons were all conducted on a Windows computer with an AMD Ryzen 7 4700U 2.00 GHz processor, featuring 8 cores and 16 GB of RAM.
Figure 22: Plot showing computation time results as features increase, faceted by observations. The points and error bars represent the minimum, median, and maximum times from 12 Monte Carlo simulations.
Figure 23: Plot showing computation time results as the number of observations increase, faceted by variables. The points and error bars represent the minimum, median, and maximum times from 12 Monte Carlo simulations.
Figure 22 shows the computational time in seconds for all three methods as a function of the number of features and for different sample sizes. From these results, we observe that both SHAP and CLIQUE are significantly faster than LIME. Additionally, the figure reveals that the SHAP runtime is approximately independent of the number of features, although the variance can be high. In contrast, the CLIQUE runtime increases linearly with the number of features and is more stable.
Figure 23 shows the same results with the computational time as a function of sample size across different numbers of features. From these results, we observe that the SHAP runtime generally increases quadratically with the number of observations, while the CLIQUE runtime remains linear.
These results suggest that CLIQUE scales better with increasing observations, while SHAP scales better with more features. This aligns with the expectations set by the authors of TreeSHAP [18], who describe a fast SHAP implementation with time complexity $O(TLD^{2})$, where $T$ is a model hyperparameter, $L$ is the maximum number of tree nodes (which is largely dependent on the number of observations $n$), and $D=log(L)$. Notably, the SHAP time complexity does not depend on the number of features $p$, implying that computational cost remains roughly constant across different values of $p$. While $p$ does influence $L$, this relationship is not monotonic and is significantly weaker than the effect of $n$.